I am trying to check my work - help
2x + 4y =50, 3x + 5 y=66
2x + 4y = 50
2x = 50 - 4y
x = 25 -2y
3(25 - 2y) + 5y = 66
75 -6y +5y =66
-y = 66 - 75 = -9, y=9
2x + 36 = 50
2x = 50 - 36 =14
x=7

Answer by deborabr(181)

i looked your work and it is perfect! congratulations!
2x + 4y =50, 3x + 5 y=66
2x + 4y = 50
2x = 50 - 4y
x = 25 -2y
3(25 - 2y) + 5y = 66
75 -6y +5y =66
-y = 66 - 75 = -9, y=9
2x + 36 = 50
2x = 50 - 36 =14
x=7